OpenStudy (anonymous):

What is M? dM/dt=(100-kM) k is a constant

jimthompson5910 (jim_thompson5910):

\[\Large \frac{dM}{dt} = 100 - kM\] \[\Large dM = (100 - kM)*dt\] \[\Large \frac{dM}{100 - kM} = dt\] \[\Large \int\frac{dM}{100 - kM} = \int 1*dt\] I'll let you finish

OpenStudy (anonymous):

I got to that part. I don't know how to integrate the left side. How do I take care of the kM. I would have to multiply by -1/M right?

jimthompson5910 (jim_thompson5910):

let u = 100 - kM and then derive both sides with respect to M u = 100 - kM du/dM = -k and then isolate dM du/dM = -k du = -k*dM -du/k = dM dM = -du/k

jimthompson5910 (jim_thompson5910):

\[\Large \int\frac{dM}{100 - kM} = \int 1dt\] turns into \[\Large \int \frac{-\frac{du}{k}}{u} = \int 1dt \] \[\Large -\frac{1}{k}\int \frac{du}{u} = \int 1dt \] \[\Large -\frac{1}{k}\int \frac{1}{u}du = \int 1dt \]

OpenStudy (anonymous):

so is the answer -1/k*ln|100-kM|=t

OpenStudy (anonymous):

oh + C

jimthompson5910 (jim_thompson5910):

yeah \[\Large -\frac{1}{k}ln(|100-kM|) = t+C\] from here you solve for M

OpenStudy (anonymous):

How do I take care of the absolute value sign?

jimthompson5910 (jim_thompson5910):

if |x| = k, then x = k or x = -k for some positive number k put another way, if |x| = k, then x = +-k

OpenStudy (anonymous):

so I have to put +- in front of the right side?

jimthompson5910 (jim_thompson5910):

Like this \[\Large -\frac{1}{k}ln(|100-kM|) = t+C\] \[\Large ln(|100-kM|) = -k(t+C)\] \[\Large |100-kM| = e^{-k(t+C)}\] \[\Large 100-kM = \pm e^{-k(t+C)}\]

OpenStudy (anonymous):

Okay I think I understand now. What a struggle. Thanks